Total no. of solution for


    2^x + 3^x + 4^x - 5^x   is


1 /  2/ 3/  or what


 

Grade:

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question


f(x) = 2^x + 3^x + 4^x - 5^x
For x value less than zero 5xwill lay below
2x+ 3x+ 4x
So there would be no solution less than zero.
Lets try greater than zero.
f(1) = 2 + 3 + 4 – 5 = 4
f(2) = 4 + 9 + 16 – 25 = 4
f(3) = 8 + 27 + 64 – 125 = -26
So 2 & 3 have opposite sign , there would be one solution b/w 2 & 3.